What is the degree measure for 3/5 rotation counterclockwise in a circle
Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

What is the degree measure for 3/5 rotation counterclockwise in a circle
----
(3/5)360 = 216 degrees
========================
Cheers,
Stan H.
=============

Answer by jim_thompson5910(35256)

(3/5)*(360) = (3*360)/5 = 1080/5 = 216


So the 3/5 of a rotation is 216 degrees.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim